OpenStudy (anonymous):

Find the difference. (5x2 + 2x + 11) - (7 + 4x - 2x2)

OpenStudy (anonymous):

the (5x2 part is an xponet so is the 2x2

OpenStudy (anonymous):

dont know how to do this would really appreciate it if someone would help me

OpenStudy (anonymous):

\[Find the difference. (5x ^2 + 2x + 11) - (7 + 4x - 2x ^2)\]

OpenStudy (anonymous):

distribute the -1 on -1(7+4x -2x^2) and then combine like terms

OpenStudy (anonymous):

could you give me an idea on what the answer should look like so i can get a better idea on how to work them please and thankx

OpenStudy (anonymous):

did you distibute the -1? what did you get?

OpenStudy (anonymous):

well i dont really understand distrubuting :/ i usually learn wheni see an example of the problem done because i have 4 consectutive problems after this one

OpenStudy (anonymous):

that invoivle disturbuting

OpenStudy (anonymous):

well difference :/ my bad

OpenStudy (anonymous):

k let's test that here's a problem: (6x^2 + 12x + 1) - (9 - 3x - 4x^2) distrubting: -1*9 + -1*-3x + -1*-4x^2 = -9 +3x +4x^2

OpenStudy (anonymous):

so algebraic you dont really have to do anything to the (6x^2 + 12x + 1)

OpenStudy (anonymous):

combining like terms: 6x^2 + 12x + 1 -9 +3x +4x^2 = 4x^2 + 6x^2 +12x +3x +1 -9 = 10x^2 +15x -8

OpenStudy (anonymous):

o i see you do

OpenStudy (anonymous):

not helpful

OpenStudy (anonymous):